OpenStudy (anonymous):

Solve the following system of equations: x − 2y = 14 x + 3y = 9

OpenStudy (anonymous):

(1, 12) (−1, −12) (12, −1) (12, 1)

OpenStudy (anonymous):

I am learning how to solve systems of equations by substitution

OpenStudy (anonymous):

Btw its actually supposed to look like this: Solve the following system of equations: {x − 2y = 14 {x + 3y = 9

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

x-2y=14 x=14+2y now substitute for x= 14+2y in the second equation 14+x=9 14+2y+3y=9 14+5y=9 5y=9-14 5y=-5 y=-1 now substitute the value of y=-1 in the first equation i.e. x-2y=14 x-2(-1)=14 x+2=14 x=12

OpenStudy (nurali):

x − 2y = 14_________(1) x + 3y = 9________(2) in equation (1) x-2y=14 x=14+2y put 'x' in equation (2) x + 3y = 9 (14+2y)+3y=9 14+2y+3y=9 5y=9-14 5y=-5 y=-1 put 'y' in equation (1) x − 2y = 14 x − 2(-1) = 14 x+2=14 x=12 (x,y)=(12,-1)

OpenStudy (anonymous):

No she did good! I actually understand it, she wrote it all out and I was able to see how she did it!

OpenStudy (anonymous):

@susan3 and @Nurali Thanks so much! You both showed me how to do it wonderfully. I understand a lot better now!

OpenStudy (nurali):

My Pleasure.